OpenStudy (anonymous):

Find the 7th term of a geometric sequence with t1=5 and r = -2.

OpenStudy (anonymous):

one of these 300 320 340 360

OpenStudy (solomonzelman):

\(\Large\color{slate}{\displaystyle~{\rm t}_{n}={\rm t}_{1}\times \left({\rm r}\right)^{n-1} }\)

OpenStudy (solomonzelman):

\(\Large\color{slate}{\displaystyle~{\rm t}_{7}=5\times \left({\rm -2}\right)^{7-1} }\)